The circumference of the exceeds it's radius by 111cm, find radius of the circle
Answer by Alan3354(69443)

The circumference of the ?? circle? exceeds it's [sic] radius by 111cm, find radius of the circle.
-----
C = 2pi*r = r + 111
2pi*r - r = 111
r*(2pi - 1) = 111
r = 111/(2pi - 1)
r =~ 21.01 cm
====================
it's = it is